A man has two kinds of money, dimes and half dimes. If he is offered $1.35 for 20 coins, how many of each kind must he give?
Answer by mananth(16949)

dimes x
nickels y
Total 750
1 x + 1 y = 20 .............1
10 x + 5 y = 135 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 1
-5 x -5 y = -100
10 x + 5 y = 135
Add the two equations
5 x = 35.00
/ 5
x = 7
plug value of x in (1)
1 x + 1 y = 20
7 + 1 y = 20
1 y = 20 -7
1 y = 13
y = 13